OpenStudy (anonymous):

Could someone please show the working for this simultaneous equation 280=x*0.177^y , 324=x*0.259^y x=543.862, y=0.383405

OpenStudy (anonymous):

280=x*0.177^y -(1) 324=x*0.259^y -(2) (2) / (1) \[\frac{0.259^y}{0.177^y} = \frac{324}{280}\]\[(\frac{0.259}{0.177})^y = \frac{324}{280}\] Then, take log on both side, and you can find y.